# Self-Consistency Preference Optimization

### v2

@@ 4.2 ZebraLogic: A Challenging Logical Reasoning Task
Table 3 reports performance on ZebraLogic of ScPO and various baselines, using greedy decoding. We observe large improvements over the seed model, Llama-3 Instruct 8B ($M_{0}$) with one iteration of unsupervised ScPO ($M_{1}$), improving performance by $5.4\%$ and $8.5\%$ in overall puzzle accuracy (exact match of tables) and cell accuracy (match of each cell in the table), respectively. In contrast, unsupervised training of $\text{IRPO}_{\mathrm{RM}}$ yields only mild gains over the seed model by $3\%$ in cell accuracy and even a slight drop in puzzle accuracy ($11.6\%$ to $11.3\%$). This can be attributed to ZebraLogic puzzles being out-of-distribution for the ArmoRM (cf. Section 5), thus trailing behind one iteration of ScPO by $5.7\%$ in puzzle accuracy and $5.5\%$ in cell accuracy. Overall, training with ScPO for two iterations improves the performance of the seed model by 8 positions on the leaderboard (from $38^{\mathrm{th}}$ to $30^{\mathrm{th}}$) with a $6.5\%$ boost in puzzle accuracy and, to the best of our knowledge, is the best 8B-scale LLM on ZebraLogic.
8B LLM trained with ScPO outperforms larger models.
Comparison of ScPO-trained models to other models in Table 3 demonstrates that ScPO-training after two iterations $(M_{2})$ outperforms significantly larger models such as Llama-3 Instruct 70B, Gemma-2 27B, and Claude-3 Haiku by $0.9\%$, $1.8\%$, and $3.8\%$ in overall puzzle accuracy, respectively. Additionally, we find that models trained using ScPO also yield the highest cell accuracy. We attribute these gains over larger models to the substantial improvement in solving easy puzzles with ScPO (up to $10.3\%$).
@@ 5 Ablations and Analysis
Figure 3: Comparing the quality of metrics: self-consistency (SC) and ArmoRM to distinguish between correct and incorrect responses.
Our results in Section 4 show that models trained with unsupervised ScPO outperform models trained with IRPO using ArmoRM to build preference pairs. To study this further, we conduct additional analysis by measuring the ability of the two methods to distinguish between correct and incorrect responses, comparing the methods to gold labels. Results are given in Figure 3. We find that ArmoRM consistently has more incorrect orderings of pairwise preferences (the chosen is incorrect and the rejected is correct) than ScPO across all three datasets (shown in red). This added noise in training may be a major factor as to why IRPORM performs poorly compared to ScPOUnsup. On the other hand, self-consistency results in a greater number of ties, i.e., when the chosen and rejected answers get the same number of votes; these are ignored in ScPO’s loss since $w(x)\!=\!0$. Lastly, we find that in out-of-distribution settings like ZebraLogic, self-consistency outperforms ArmoRM with $12.3\%$ higher correct orderings of pairwise preferences (shown in green in Figure 3).
@@ 6 Related Work
Iterative training or self-training has shown meaningful improvements in a number of domains such as safety (Bai et al., 2022), multilingual reasoning (She et al., 2024), and evaluation (Wang et al., 2024b). Because LLMs often struggle with both generating and validating solutions to complex reasoning tasks, prior works on training LLMs for complex problem-solving tasks largely rely on human-annotated (gold) final answers (Zelikman et al., 2022; Chen et al., 2024b; Pang et al., 2024) or access to an external reward model that performs well on the underlying task (Singh et al., 2024; Dong et al., 2023). However, both these classes of approaches suffer from their own shortcomings. Firstly, manually annotating or verifying the final answer requires working through the solution step-by-step, making it especially resource-intensive for complex multi-step problems. Training strong reward models for such reasoning and problem-solving tasks also often requires human judgements of LLM generations (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2023), making it similarly expensive. Our work focuses on the setting without access to gold solutions or final answers, which remains largely unaddressed. While other works such as She et al. (2024); Yuan et al. (2024); Rosset et al. (2024); Tran et al. (2023) geared towards general instruction following tasks (as opposed to reasoning tasks specifically) circumvent the need for human-annotated labels in the dataset by using the model itself to score the responses, these works demonstrate only modest gains in the context of reasoning tasks.
@@
Self-consistency (Wang et al., 2023) relies upon the intuition that sampling several responses, some of which lead to the same answer, lends higher certainty that the consistent answer is the correct one. Application of self-consistency at inference time has enabled performance improvements in a number of domains like math (Wang et al., 2023), code generation (Shi et al., 2022; Li et al., 2022; Chen et al., 2018), and even open-ended tasks like summarization and question answering (Chen et al., 2024a). In this work, we explore using self-consistency at training time for reasoning tasks, constructing preference pairs according to the self-consistent final answer. While Huang et al. (2023) also use self-consistency to finetune models without access to gold labels via NLL loss, we employ a preference optimization loss function that is weighted according to the consistency of an answer. Intuitively, the consistency of an answer is a reflection of the model confidence, and several prior works have demonstrated that leveraging model uncertainty can lead to faster convergence and improved performance (Gal & Ghahramani, 2016; Krishnan & Tickoo, 2020; Corbière et al., 2019).
@@ 7 Conclusion
In this paper, we introduced Self-Consistency Preference Optimization (ScPO). ScPO leverages the concept of self-consistency, usually employed only at inference time, to improve the self-training of large language models. By iteratively optimizing to prefer consistent answers to inconsistent ones, ScPO achieves significant improvements over traditional reward model training without the need for additional gold labels. Our experiments demonstrate the efficacy of ScPO on various reasoning tasks, including GSM8K, MATH, and ZebraLogic, where in the latter it outperforms several larger state-of-the-art language models. We also showed that ScPO works well in semi-supervised setups with access to some gold labels, in addition to unlabeled inputs – improving performance further. These results highlight the potential of ScPO to improve self-alignment across reasoning tasks – a domain that prior self-alignment methods still struggle with. Future work could extend ScPO to tasks where a single final answer cannot be easily parsed (e.g., summarization) through universal self-consistency (Chen et al., 2024a), which leverages an LLM to select the most consistent answer among multiple samples. While we explore consistency in this work according to one model class (Llama-3 8B Base and Instruct), future work could also investigate consistency according to a suite of other models and tasks.
@@ Appendix A Relationship between Consistency and Accuracy
Level of consistency or vote share correlates with accuracy.
@@
We observe that the degree of consistency, or vote share, is positively and strongly correlated with accuracy. This relationship is evidenced in Table 6 by a high rank order correlation for all three datasets, as determined by Somer’s D (Somers, 1962), which measures the degree of association between two possibly dependent variables. This association is lowest for MATH, likely because the challenging nature of this task makes it difficult for the model to produce consistent answers.
@@
Table 7: Additional iteration of training ($M_{3}$) by bootrapping from questions in the train and test set. On GSM8K, we bootstrap 8.7K pairs using train problems and 5.8K pairs with the test problems. On MATH, we build 4.4K, and 4.2K preference pairs using train and test problems, respectively.
Appendix B Transduction During Inference
Bootstrapping preference pairs from test queries further boosts performance.
In our primary experiments, we report results for two rounds of iterative training. However, as shown in Table 7, introducing a third round of ScPO yields only marginal improvements, with gains of less than 1% over the second round. To address this saturation, we explore generating new problems and building preference pairs using the queries from test split as exemplars instead of the train split. This strategy results in more substantial improvements (+1.44% for GSM8K), as it enables the model to better adapt to the unique characteristics of the test set. For MATH, we see more substantial improvements when using SC accuracy, resulting in an improvement bump of 1.26%. We note that ZebraLogic is excluded from this analysis, as it only provides test samples.
Appendix C Prompts
In this section, we provide all task-specific prompts used for both generating new problems and for generating candidate solutions.

### v3

@@ 4.2 ZebraLogic: A Challenging Logical Reasoning Task
Table 3 reports performance on ZebraLogic of ScPO and various baselines, using greedy decoding. We observe large improvements over the seed model, Llama-3 Instruct 8B ($M_{0}$) with one iteration of unsupervised ScPO ($M_{1}$), improving performance by $5.4\%$ and $8.5\%$ in overall puzzle accuracy (exact match of tables) and cell accuracy (match of each cell in the table), respectively. In contrast, unsupervised training of $\text{IRPO}_{\mathrm{RM}}$ yields only mild gains over the seed model by $3\%$ in cell accuracy and even a slight drop in puzzle accuracy ($11.6\%$ to $11.3\%$). This can be attributed to ZebraLogic puzzles being out-of-distribution for the ArmoRM (cf. Section 5), thus trailing behind one iteration of ScPO by $5.7\%$ in puzzle accuracy and $5.5\%$ in cell accuracy. Moreover, two iterations of ScPO outperform that of LMSI by $4.6\%$ on easy puzzles and $1.3\%$ on overall accuracy. Taken together, training with ScPO for two iterations improves the performance of the seed model by 8 positions on the leaderboard (from $38^{\mathrm{th}}$ to $30^{\mathrm{th}}$) with a $6.5\%$ boost in puzzle accuracy and, to the best of our knowledge, is the best 8B-scale LLM on ZebraLogic.
8B LLM trained with ScPO outperforms larger models. Comparison of ScPO-trained models to other models in Table 3 demonstrates that ScPO-training after two iterations $(M_{2})$ outperforms significantly larger models such as Llama-3 Instruct 70B, Gemma-2 27B, and Claude-3 Haiku by $0.9\%$, $1.8\%$, and $3.8\%$ in overall puzzle accuracy, respectively. Additionally, we find that models trained using ScPO also yield the highest cell accuracy. We attribute these gains over larger models to the substantial improvement in solving easy puzzles with ScPO (up to $10.3\%$).
@@ 5 Ablations and Analysis
Figure 3: Comparing the quality of metrics: self-consistency (SC) and ArmoRM to distinguish between correct and incorrect responses on all datasets.
Our results in Section 4 show that models trained with unsupervised ScPO outperform models trained with IRPO using ArmoRM to build preference pairs. To study this further, we conduct additional analysis by measuring the ability of the two methods to distinguish between correct and incorrect responses, comparing the methods to gold labels in Figure 3. We find that ArmoRM consistently has more incorrect orderings of pairwise preferences (the chosen is incorrect and the rejected is correct) than ScPO across all three datasets (shown in red). This added noise in training may be a major factor as to why IRPORM performs poorly compared to ScPOUnsup. On the other hand, self-consistency results in a greater number of ties, i.e., when the chosen and rejected answers get the same number of votes; these are ignored in ScPO’s loss since $w(x)\!=\!0$. Lastly, we find in the out-of-distribution setting of ZebraLogic, self-consistency outperforms ArmoRM with $12.3\%$ more correct orderings of pairwise preferences (shown in green in Figure 3).
@@ 6 Related Work
Iterative training or self-training has shown meaningful improvements in a number of domains such as safety (Bai et al., 2022), multilingual reasoning (She et al., 2024), and evaluation (Wang et al., 2024b). Because LLMs often struggle with both generating and validating solutions to complex reasoning tasks, prior works on training LLMs for complex problem-solving tasks largely rely on human-annotated (gold) final answers (Zelikman et al., 2022; Chen et al., 2024b; Pang et al., 2024) or access to an external reward model that performs well on the underlying task (Singh et al., 2024; Dong et al., 2023). However, both these classes of approaches suffer from their own shortcomings. Firstly, manually annotating or verifying the final answer requires working through the solution step-by-step, making it especially resource-intensive for complex multi-step problems. Training strong reward models for such reasoning and problem-solving tasks also often requires human judgements of LLM generations (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2024), making it similarly expensive. Our work focuses on the setting without access to gold solutions or final answers, which remains largely unaddressed. While other works such as She et al. (2024); Yuan et al. (2024); Rosset et al. (2024); Tran et al. (2023) geared towards general instruction following tasks (as opposed to reasoning tasks specifically) circumvent the need for human-annotated labels in the dataset by using the model itself to score the responses, these works demonstrate only modest gains in the context of reasoning tasks.
@@
Self-consistency (Wang et al., 2023) relies upon the intuition that sampling several responses, some of which lead to the same answer, lends higher certainty that the consistent answer is the correct one. Application of self-consistency at inference time has enabled performance improvements in a number of domains like math (Wang et al., 2023), code generation (Shi et al., 2022; Li et al., 2022; Chen et al., 2018), and even open-ended tasks like summarization and question answering (Chen et al., 2024a). In this work, we explore using self-consistency at training time for reasoning tasks, constructing preference pairs according to the self-consistent final answer. While Huang et al. (2023) also use self-consistency to finetune models without access to gold labels via NLL loss, we employ a preference optimization loss function that is weighted according to the consistency of an answer. Intuitively, the consistency of an answer is a reflection of the model confidence, and several prior works have demonstrated that leveraging model uncertainty can lead to faster convergence and improved performance (Gal & Ghahramani, 2016; Krishnan & Tickoo, 2020; Corbière et al., 2019). Concurrently with this work, Jiao et al. (2025) propose training models on “pseudo-feedback” from test cases, wherein they employ self-consistency to construct the test cases itself. However, we note that our work additionally shows the utility of self-consistency in generating new problems to augment the seed data (Section 4) as well as in our weighted loss function (Table 4 in Section 5).
@@ 7 Conclusion
In this paper, we introduced Self-Consistency Preference Optimization (ScPO). ScPO leverages the concept of self-consistency, usually employed only at inference time, to improve the self-training of large language models. By iteratively optimizing to prefer consistent answers to inconsistent ones, ScPO achieves significant improvements over traditional reward model training without the need for additional gold labels. Our experiments demonstrate the efficacy of ScPO on various reasoning tasks, including GSM8K, MATH, and ZebraLogic, where in the latter it outperforms several larger state-of-the-art language models. We also showed that ScPO works well in semi-supervised setups with access to some gold labels, in addition to unlabeled inputs – improving performance further. These results highlight the potential of ScPO to improve self-alignment across reasoning tasks – a domain that prior self-alignment methods still struggle with. Future work could extend ScPO to tasks where a single final answer cannot be easily parsed (e.g., summarization) through universal self-consistency (Chen et al., 2024a). While we explore consistency according to several models (Llama-3 and 3.1 8B, Base and Instruct), future work could also investigate consistency according to a suite of other models and tasks.
Acknowledgements
We sincerely thank Ilia Kulikov, other members of the RAM team at FAIR, as well as the anonymous reviewers for their valuable feedback on the paper. Part of this work was done during an internship at Meta FAIR and was partially supported at UNC by NSF-CAREER Award 1846185, NSF-AI Engage Institute DRL-2112635, DARPA Machine Commonsense (MCS) Grant N66001-19-2-4031. The views contained in this article are those of the authors and not of the funding agencies.
Impact Statement
This work presents a new training algorithm that uses self-consistency for training large language models on math and logical reasoning tasks without the need for gold labels. The outputs produced by models trained with ScPO may exhibit undesirable behavior similar to the base model and have the same potential for misuse as other fine-tuned LLMs (Weidinger et al., 2021). Hence, more studies are needed to evaluate and mitigate such biases in LLMs.
@@ Appendix A Relationship between Consistency and Accuracy
Level of consistency or vote share correlates with accuracy. We observe that the degree of consistency, or vote share, is positively and strongly correlated with accuracy. This relationship is evidenced in Table 6 by a high rank order correlation for all three datasets, as determined by Somer’s D (Somers, 1962), which measures the degree of association between two possibly dependent variables. This association is lowest for MATH, likely because the challenging nature of this task makes it difficult for the model to produce consistent answers.
@@
Table 7: Somers’ D computed between $\mathrm{Acc}(y)$ and $\mathcal{V}(y)$ for $y\in\{y^{+},y^{-}\}$, i.e., the most and least consistent responses, on test set for different values of $k$.
Dataset / Somer’s D $k=2$ $k=4$ $k=8$ $k=16$
GSM-8K 0.39 0.65 0.80 0.89
ZebraLogic 0.66 0.82 0.92 0.93
Furthermore, we measure the impact of the number of samples used to measure self-consistency ($k$) on its Somer’s D correlation with correctness in Table 7. The results indicate that (i) lower values of $k$ (e.g. $k=2/4$) have lower correlation with correctness or accuracy which we find is because of fewer instances where any answer gets multiple votes; (ii) while larger values of $k=16$ yield slightly higher correlations, we prioritize computational efficiency in the data generation phase, and use a sufficiently large value of $k=8$ in addition to filtering and a weighted loss in ScPO.
Table 8: GSM8K zero-shot accuracy after training Llama-3.1 Base 8B with ScPO and baselines, using greedy or self-consistency (SC)-based inference.
{NiceTabular}
llr@ / lcc Method Iter. Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC (8-way)
[…]
 Seed model (zero-shot) $M_{0}$ 43.14 59.59
 IRPORM $M_{1}$ 6.5 - 58.60 73.01
$M_{2}$ 6.7 - 60.04 72.19
 LMSI $M_{1}$ 6.7 5.7 48.75 65.71
$M_{2}$ 6.3 4.8 52.39 60.42
 ScPOUnsup. $M_{1}$ 6.7 5.7 61.64 71.95
$M_{2}$ 5.5 4.9 64.22 75.13
[…]
 IRPOGold $M_{1}$ 5.6 - 60.05 76.04
$M_{2}$ 5.8 - 65.50 79.61
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ $M_{1}$ 5.6 5.4 65.60 79.08
$M_{2}$ 5.2 4.9 68.46 79.75
Table 9: MATH zero-shot accuracy after training Llama-3.1 Base 8B with ScPO and baselines, using greedy or self-consistency (SC)-based inference.
{NiceTabular}
llr@ / lcc Method Iter., Train Data (K) Test Acc. (%)
# Seed Gen. Greedy SC (8-way)
[…]
 Seed model (zero-shot) $M_{0}$ 15.70 24.62
 IRPORM $M_{1}$ 6.2 - 20.68 27.32
$M_{2}$ 6.6 - 20.74 25.88
 LMSI $M_{1}$ 0.9 0.9 16.26 24.38
$M_{2}$ 1.0 1.3 15.94 22.60
 ScPOUnsup. $M_{1}$ 0.9 0.9 19.38 27.74
$M_{2}$ 1.4 1.7 23.20 30.10
[…]
 IRPOGold $M_{1}$ 2.7 - 22.40 31.64
$M_{2}$ 3.2 - 22.86 32.30
 ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ $M_{1}$ 2.7 0.9 22.98 32.18
$M_{2}$ 3.2 2.2 24.36 32.64
Appendix B Transduction During Inference
Bootstrapping preference pairs from test queries further boosts performance. In our primary experiments, we report results for two rounds of iterative training. However, as shown in Table 10, introducing a third round of ScPO yields only marginal improvements, with gains of less than 1% over the second round. To address this saturation, we explore generating new problems and building preference pairs using the queries from test split as exemplars instead of the train split. This strategy results in more substantial improvements (+1.44% for GSM8K), as it enables the model to better adapt to the unique characteristics of the test set. For MATH, we see more substantial improvements when using SC accuracy, resulting in an improvement bump of 1.26%. We note that ZebraLogic is excluded from this analysis, as it only provides test samples.
Appendix C Results on Math Reasoning with Llama-3.1
We now repeat the math reasoning experiments in Section 4.1 with Llama-3.1 Base 8B and find that while the absolute performance increases, the relative trends among the baselines remain the same – with ScPOUnsup. as the most performant unsupervised technique and ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ yielding the overall highest accuracy on GSM8K and MATH. In Appendices A and A, we observe that two iterations of ScPOSemi-${}_{\mathrm{Semi\text{-}Sup.}}$ improve the greedy test accuracy of the seed model by $25.32\%$ and $8.66\%$ on GSM8K and MATH, respectively; while two iterations of ScPOUnsup. boost the greedy accuracy of the seed model by $21.08\%$ on GSM8K and $7.5\%$ on MATH dataset.
Table 10: Training $M_{3}$ by bootrapping from questions in the train and test set. On GSM8K, we bootstrap 8.7K, 5.8K pairs using train, and test problems, respectively. On MATH, we build 4.4K, and 4.2K preference pairs using train and test problems, respectively.
@@ Appendix B Transduction During Inference
Appendix D Prompts
We provide all task-specific prompts used for both generating new problems and for generating candidate solutions.